* If the price of coffee (Good 2) rises from Rs. 45 per to Rs. 55 per pack and as a result the consumer’s demand for tea (Good 1) increases from 600 to 800 packs then calculate cross elasticity of demand of tea for coffee


1
Expert's answer
2022-03-27T18:50:07-0400

Cross Elasticity of Demand ="\frac{ increase.in. quantity. of. tea}{ increase .in. the. price. of .coffee}"

"= \frac{(800-600)}{800+600} / \frac{(55-45)}{(55+45)} = 1.4285"